### v1

@@ 4.3 Lorenz’63 System
Lorenz63 is a classic, chaotic system model proposed by American meteorologist Edward Lorenz in 1963. It describes the nonlinear dynamical behavior in atmospheric circulation. The Lorenz63 model is based on three partial differential equations and has three state variables: $X_{1}$, $X_{2}$, and $X_{3}$. This model is known for its highly sensitive initial conditions and nonlinear coupling. It leads to the famous “butterfly effect” where a small perturbation of initial condtion may result in long-term uncertainty and unpredictable outcomes. It can expressed in the form of ODEs
(35) $\frac{\mathrm{d}X_{1}}{\mathrm{d}t}$ $=-\sigma X_{1}+\sigma X_{2}+\gamma_{1}\frac{\mathrm{d}B_{1}}{\mathrm{d}t},$
@@
where $B_{1},B_{2}$ and $B_{3}$ are standard Brownian motion. Here, the parameters are set to be $\beta=8/3,\rho=28,\sigma=10$, which produces the well-known Lorenz attractor. The noise level of the different components is set to be the same $\gamma_{1}=\gamma_{2}=\gamma_{3}=1.0$. We simulate the ODEs Eq. 35 by using a fourth-order Runge-Kutta method with time stepsize $\Delta t=0.01$. Here, we only take the observation operator to observe its first component. Partial observations greatly enhance the complexity and nonlinearity of the filtering problem. At the same time, we consider observations with Gaussian white noise.
@@
At first, we assimilate the Lorenz’63 system with the default setup of the ensemble transport filter. Fig. 9 and Fig. 10 respectively show the average RMSE and Spread versus a range of observation interval and ensemble size.
@@
Figure 9: Average RMSE of Lorenz’63 system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
As shown in the left panel of Fig. 9, we demonstrate the average RMSE of EnTranF as a function of observation interval with a fixed ensemble size. EnTranF(L-L) and EnTranF(N-L) experienced a collapse in the Lorenz’63 system, rendering it unable to produce meaningful state estimates. This is because the linear kernel function results in significant information loss at each assimilation window, making the method ineffective as sequential filtering progresses. EnTranF with a Gaussian kernel still performs well. EnTranF(L-G) and EnTranF(N-G) show a slight performance improvement compared to EnKF when the observation interval is relatively short. However, for sufficiently large observation intervals, they present an increasing improvement on RMSE compared to EnKF as the observation interval increases. EnTranF(N-G) and EnTranF(L-G) have $25.30\%$ and $18.63\%$ improvement on RMSE, respectively, when taking observation interval as $\Delta t_{ob}=0.5$. A well-chosen kernel function can make EnTranF more robust in systems with strong nonlinearity. The right panel of Fig. 9 shows the average RMSE of EnTranF as a function of ensemble size for the observation interval to be fixed. Since the EnTranF is based on Particle Filter, it exhibits almost no advantage over EnKF and may even perform worse when the ensemble size is small ($N<20$). However, EnTranF performs better than EnKF in the filtering problem of the Lorenz’63 system at larger aggregate ensemble sizes. If ensemble size $N\geq 100$, EnTranF(L-G) has about $18.56\%$ improvement on average RMSE compared to EnKF, and EnTranF(N-G) has about $25.04\%$ improvement. Unlike the one-dimensional double-well problem, the EnTranF with a nonlinear transport map has apparent advantages over the linear version. The relatively higher dimensional lorenz’63 system exhibits strong nonlinearities in partial observations, making methods based on linear transport maps less effective.
@@
Figure 10: Spread of Lorenz’63 system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
The ensemble spread of different filtering methods is shown in Fig. 10. EnTranF(N-G) reduces the ensemble spread compared to EnTranF(L-G). However, both EnTranF(N-G) and EnTranF(L-G) have a larger ensemble spread with different observation intervals and ensemble sizes. This may be caused by inadequate optimization of the MMD loss function. EnTranF fails to approximate the posterior adequately. The EnTranF with a variance penalty can significantly improve the spread and slightly improve the average RMSE. Fig. 11 and Fig. 12 respectively depict the average RMSE and Spread versus a range of observation intervals and ensemble sizes.
@@
Figure 11: EnTranF with variance penalty: Average RMSE of Lorenz’63 system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
Figure 12: EnTranF with variance penalty: Spread of Lorenz’63 system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
The average RMSE of EnTranF with a variance penalty outperforms EnKF at different observation intervals (see left panel of Fig. 11). After introducing the variance penalty term, even EnTranFp with a linear kernel function shows a significant performance improvement on average RMSE compared to EnKF. EnTranF(L-L) has $14.20\%$ improvement, and EnTranF(N-L) has $23.57\%$ improvement. EnTranFp also shows some improvement compared to EnTranF when taking a Gaussian kernel function. Here, EnTranFp(L-G) is $21.07\%$ better than EnKF on RMSE at observation interval $\Delta t_{ob}=0.5$, while EnTranFp(N-G) is $27.39\%$ better than EnKF. From panel (b) of Fig. 11, EnTranFp has a smaller average RMSE when $N\geq 20$ than EnKF for different ensemble sizes. In Fig. 12, EnTranFp(L-L), EnTranFp(N-L), and EnTranFp(L-G) show a similar spread to EnKF but get a higher probability of covering the true state. EnTranFp(N-G) performs optimally in the Lorenz63 system. It significantly reduces the spread while maintaining a high probability of covering the true state. The variance penalty term effectively guides EnTranF in prioritizing the approximation of high-informative statistics of the posterior distribution. It enhances the robustness of EnTranF in high-dimensional nonlinear systems.
@@ 4.4 Lorenz’96 System
The Lorenz96 model is a simplified mathematical model that describes the nonlinear dynamical behavior in atmospheric circulation systems. It consists of a set of coupled one-dimensional differential equations, each representing a grid point within the atmospheric circulation system. These differential equations are typically given by:
(36) $\frac{{\mathrm{d}X_{i}}}{{\mathrm{d}t}}=(X_{i+1}-X_{i-2})X_{i-1}-X_{i}+F.$
@@
Here, $X_{i}$ represents the state variable at the $i$-th grid point, and $F$ is an external forcing.
The numerical example is conducted to evaluate the performance of EnTranF for a of Lorenz’96 system in which sparse observations are given both in space and time. We take 20-dimension state variables and observe 10 dimensions of them (observing every other state component). We simulate the ODEs Eq. 36 by using a fourth-order Runge-Kutta method with time stepsize $\Delta t=0.01$. The observation interval in time is set to be $\Delta_{ob}=0.1$. Results of EnTranF(N-G) are presented in Fig. 13 and Fig. 14(b). From Fig. 13, EnTranF(N-G) successfully achieves the estimation of the Lorenz96 state, with error performance similar to that of EnKF. Panel (a) and Panel (b) of Fig. 14(b) respectively present the RMSE and Spread of EnTranF(N-G) and EnKF. We note that EnTranF(N-G) is comparable to EnKF in RMSE and the spread.
@@
Figure 13: State estimation and absolute error of Lorenz’96 system for observation interval $\Delta t_{ob}=0.1$ and ensemble size $N=2000$.
(a) RMSE
(b) Ensemble Spread
@@
Figure 14: RMSE and Spread of Lorenz96 system for observation interval $\Delta t_{ob}=0.1$ and ensemble size $N=2000$.
@@ 5 Conlusion
In this paper, we reconstructed the analysis step of the particle filter through a transport map, which directly transported the prior particles to the posterior particles. By matching the expectation information of posterior distribution, the problem was converted into an optimization problem for the Maximum Mean Discrepancy loss function. In the ensemble transport filter, the particle weights in the analysis step of the particle filter were substituted with the update of particle states, effectively alleviating the risk of particle degeneracy in high-dimensional problems. Meanwhile, the ensemble transport filter inherited the accurate estimation of the posterior distribution from the particle filter. To improve the robustness of MMD, we introduced a variance penalty term to guide the prioritized optimization of high-informative statistics in the approximate posterior and reference posterior.
Numerical results demonstrated that the ensemble transport filter outperformed EnKF significantly in lower-dimensional problems, particularly in strongly nonlinear systems. The ensemble transport filter remained stable in high-dimensional systems and performed comparably to EnKF. The variance penalty term effectively enhanced the robustness of our method in approximating the filtering posterior, making it more effective and stable. EnTranF with a variance penalty term significantly reduced the spread of ensemble methods while maintaining RMSE.
The general transport map structure can be explored for future work to avoid computationally expensive optimization problems when constructing the transport map at each assimilation window.

### v2

@@ 4.3 Lorenz’63 System
Lorenz63 is a classic, chaotic system model proposed by American meteorologist Edward Lorenz in 1963. It describes the nonlinear dynamical behavior in atmospheric circulation. The Lorenz63 model is based on three partial differential equations and has three state variables: $X_{1}$, $X_{2}$, and $X_{3}$. This model is known for its highly sensitive initial conditions and nonlinear coupling. It leads to the famous “butterfly effect” where a small perturbation of initial condition may result in long-term uncertainty and unpredictable outcomes. It can expressed in the form of ODEs
$\frac{\mathrm{d}X_{1}}{\mathrm{d}t}$ $=-\sigma X_{1}+\sigma X_{2}+\gamma_{1}\frac{\mathrm{d}B_{1}}{\mathrm{d}t},$ (4.35)
@@
where $B_{1},B_{2}$ and $B_{3}$ are standard Brownian motion. Here, the parameters are set to be $\beta=8/3,\rho=28,\sigma=10$, which produces the well-known Lorenz attractor. The noise level of the different components is set to be the same $\gamma_{1}=\gamma_{2}=\gamma_{3}=4\times 10^{-4}$. We simulate the ODEs (4.35) by using a fourth-order Runge-Kutta method with time stepsize $\Delta t=0.01$. Here, we only take the observation operator to observe its first component. Partial observations greatly enhance the complexity and nonlinearity of the filtering problem. At the same time, we consider observations with Gaussian white noise. In this numerical example, we assimilate the Lorenz’63 system over $500$ assimilation windows. A shallow DNN comprising $L=3$ layers with width $\{1,40,3\}$ and tanh activation functions is used to define the nudging term of the nonlinear transport map. The bandwidth of MMD was set to be the median of all $L_{2}$ distance between each pair of particles.
@@
At first, we assimilate the Lorenz’63 system with the default setup of the ensemble transport filter. Figure 4.12 and Figure 4.13 respectively show the average RMSE and Spread versus a range of observation interval and ensemble size.
@@
Figure 4.12: Average RMSE of Lorenz’63 system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
As shown in the left panel of Figure 4.12, we demonstrate the average RMSE of EnTranF as a function of observation interval with a fixed ensemble size. EnTranF(L-L) and EnTranF(N-L) experienced a collapse in the Lorenz’63 system, rendering it unable to produce meaningful state estimates. This is because the linear kernel function results in significant information loss at each assimilation window, making the EnTranF ineffective in finding admissible transport maps as sequential filtering progresses. EnTranF with a Gaussian kernel still performs well. EnTranF(L-G) and EnTranF(N-G) show a similar performance compared to EnKF when the observation interval is relatively short. For sufficiently large observation intervals, all variant of EnTranF present an increasing improvement on RMSE compared to EnKF as the observation interval increases. EnTranF(N-G) and EnTranF(L-G) have $28.57\%$ and $15.89\%$ improvement on RMSE, respectively, when taking observation interval as $\Delta t_{ob}=0.5$. A well-chosen kernel function can make EnTranF more robust in systems with strong nonlinearity. The right panel of Figure 4.12 shows the average RMSE of EnTranF as a function of ensemble size for the observation interval to be fixed. Since the EnTranF is based on Particle Filter, it exhibits almost no advantage over EnKF and may even perform worse when the ensemble size is small ($N<20$). However, EnTranF performs better than EnKF in the filtering problem of the Lorenz’63 system at larger ensemble sizes. If ensemble size $N\geq 100$, EnTranF(L-G) has about $15.81\%$ improvement on average RMSE compared to EnKF, and EnTranF(N-G) has about $23.77\%$ improvement. Unlike the one-dimensional double-well problem, the EnTranF with a nonlinear transport map has apparent advantages over the linear version. The relatively higher dimensional lorenz’63 system exhibits strong nonlinearities in partial observations, making methods based on linear transport maps less effective.
@@
Figure 4.13: Spread of Lorenz’63 system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
The ensemble spread of different filtering methods is shown in Figure 4.13. EnTranF(N-G) reduces the ensemble spread compared to EnTranF(L-G). However, both EnTranF(N-G) and EnTranF(L-G) have a larger ensemble spread with different observation intervals and ensemble sizes. This may be caused by inadequate optimization of the MMD loss function. EnTranF fails to approximate the posterior adequately. The EnTranF with a variance penalty can significantly improve the spread and slightly improve the average RMSE. Figure 4.14 and Figure 4.15 respectively depict the average RMSE and Spread over a range of observation intervals and ensemble sizes.
@@
Figure 4.14: EnTranF with variance penalty: Average RMSE of Lorenz’63 system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
Figure 4.15: EnTranF with variance penalty: Spread of Lorenz’63 system for ensemble size $N=400$ as a function of observation interval (Left), and for observation interval $\Delta t_{ob}=0.5$ as a function of ensemble size (Right).
@@
The average RMSE of EnTranF with a variance penalty outperforms EnKF at different observation intervals (see left panel of Figure 4.14). After introducing the variance penalty term, even EnTranFp with a linear kernel function shows a significant performance improvement on average RMSE compared to EnKF. EnTranF(L-L) has $12.37\%$ improvement, and EnTranF(N-L) has $28.06\%$ improvement. EnTranFp also shows some improvement compared to EnTranF when taking a Gaussian kernel function. Here, EnTranFp(L-G) is $23.76\%$ better than EnKF on RMSE at observation interval $\Delta t_{ob}=0.5$, while EnTranFp(N-G) is $37.11\%$ better than EnKF. From right panel of Figure 4.14, EnTranFp has a smaller average RMSE when $N\geq 20$ than EnKF for different ensemble sizes. In Figure 4.15, EnTranFp(L-L), EnTranFp(N-L), and EnTranFp(L-G) show a similar spread to EnKF but get a higher probability of covering the true state. EnTranFp(N-G) performs optimally in the Lorenz63 system. It significantly reduces the spread while obtaining a probability of covering the true state close to $95\%$. The variance penalty term effectively guides EnTranF in prioritizing the approximation of high-informative statistics of the posterior distribution. It enhances the robustness of EnTranF in high-dimensional nonlinear systems.
@@ 4.4 Lorenz’96 System
The Lorenz’96 model is a simplified mathematical model that describes the nonlinear dynamical behavior in atmospheric circulation systems. It consists of a set of coupled one-dimensional differential equations, each representing a grid point within the atmospheric circulation system. These differential equations are typically given by:
$\frac{{\mathrm{d}X_{i}}}{{\mathrm{d}t}}=(X_{i+1}-X_{i-2})X_{i-1}-X_{i}+F+\gamma_{i}\frac{\mathrm{d}B_{i}}{\mathrm{d}t}.$ (4.36)
@@
Here, $X_{i}$ represents the state variable at the $i$-th grid point, and $F$ is an external forcing. The noise level of the different components is set to be the same $\gamma_{i}=4\times 10^{-4}$.
The numerical example is conducted to evaluate the performance of EnTranF for a of Lorenz’96 system in which sparse observations are given both in space and time. We take 40-dimension state variables and observe 20 dimensions of them (observing every other state component). We simulate the ODEs (4.36) by using a fourth-order Runge-Kutta method with time stepsize $\Delta t=0.01$. The observation interval in time was set to be $\Delta_{ob}=0.4$ and assimilation steps was set to be $500$. The observation noise is set to be white noise with a variance of $1.0$. Numerical results are presented in Figure 4.16 and Figure 4.17. To address the Lorenz’96 system, we split the 20-dimensional observation vector into components, each consisting of two dimensions. Then, prior ensembles are sequentially transform to posterior ensembles using the approach detailed in Section 3.2.2, which allows for more effective ensemble transport filter in this high-dimensional setting. For each component, the transport map was defined using a DNN comprising $L=4$ layers with widths $\{2,10,10,40\}$ and tanh activation functions to parametric the nudging term. The bandwidth of MMD was set to be the median of all $L_{2}$ distance between each pair of particles.
@@
Figure 4.16: State estimation and absolute error of EnKF and EnTranF for Lorenz’96 system with observation interval $\Delta t_{ob}=0.4$ and ensemble size $N=4000$.
Figure 4.16 shows partial states over time-horizen of Lorenz’96 estimated using EnKF and EnTranF(N-G), respectively, along with their absolute errors relative to the true state. In this 40-dimensional problem, the PF cannot be implemented with limited ensemble size due to the particle degeneracy issue. EnTranF(N-G) obtained a correct state estimation of the Lorenz’96 state, with a similar absolute error compare to EnKF. The results show that EnTranF, which reconstructs the resampling step via transport maps, is able to rescure the issue of particle degeneracy.
@@
Figure 4.17: Average RMSE of EnKF, EnTranF(L-L) and EnTranF(N-G) for Lorenz’96 system with observation interval $\Delta t_{ob}=0.4$ as a function of ensemble size.
The average RMSE of EnKF, EnTranF(L-L) and EnTranF(N-G) are shown in Figure 4.17. It presents that both EnTranF(L-L) and EnTranF(N-G) exhibit superior performance compared to the EnKF, and as the number of particles increases, the performance of EnTranF(N-G) gradually improves. When the ensemble size exceeds 1600, EnTranF achieves an average RMSE reduction of 10.34% compared to EnKF. Here, EnTranF requires a large ensemble size compare to EnKF, primarily because it relies on sampling particles to approximate these distributions and does not perform state space localization. This means that to accurately capture the characteristics of the target distribution, a large number of particles is usually required. When the ensemble size is small, the inadequacy in the number of particles can lead to insufficient accuracy in the approximation of the distribution, thereby affecting the effectiveness of the algorithm.
For smaller ensemble sizes, we compare EnTranF with LETKF both in a domain localization setting with localization radius $r_{loc}=2$. The inflation parameter for the LETKF is set to be $1.2$. In this setup, the dimension of each local block is 5 and the data assimilation for each local block can be performed in parallel. The average RMSE and coverage probability of EnKF and EnTranF are shown in Figure 4.18.
Figure 4.18: Average RMSE and Coverage probability of LETKF and EnTranF for Lorenz’96 system with observation interval $\Delta t_{ob}=0.4$ as a function of ensemble size.
From the left panel of Figure 4.18, EnTranF(N-G) achieves a lower RMSE than LETKF when the ensemble size $N$ exceeds $100$, resulting in more accurate state estimation. EnTranF(L-L) requires an ensemble size exceeding $300$ to achieve performance slightly better than that of LETKF. EnTranF(L-L) can achieve a coverage rate close to $95\%$ when the ensemble size exceeds 200, indicating accurate uncertainty estimation. EnTranF(L-L) performs generally poorly in the high-dimensional Lorenz’96 system. The linear kernel and linear map make it difficult to adequately capture the features of complex high-dimensional distributions, thereby reducing its effectiveness in sequential data assimilation tasks.
@@ 5 Conlusion
In this paper, we reconstructed the analysis step of the particle filter through a transport map, which directly transported the prior particles to the posterior particles. By matching the expectation information of posterior distribution, the problem was converted into an optimization problem for the Maximum Mean Discrepancy loss function. In the ensemble transport filter, the particle weights in the analysis step of the particle filter were substituted with the update of particle states, effectively alleviating the risk of particle degeneracy in high-dimensional problems. Meanwhile, the ensemble transport filter inherited the accurate estimation of the posterior distribution from the particle filter. To improve the robustness of MMD, we introduced a variance penalty term to guide the prioritized optimization of high-informative statistics in reference posterior.
Numerical results demonstrated that the ensemble transport filter outperformed EnKF significantly in lower-dimensional problems, particularly in strongly nonlinear systems. The ensemble transport filter remained efficient in high-dimensional systems and performed comparably to EnKF. The variance penalty term effectively enhanced the robustness of our method in approximating the filtering posterior, making it more effective and stable. EnTranF with a variance penalty term significantly reduced the spread of ensemble methods while maintaining RMSE.
The general transport map structure can be explored for future work to avoid computationally expensive optimization problems when constructing the transport map at each assimilation window. Furthermore, Motivated by the block-triangular structure of transport map [1], a likelihood-free version of our mehtod can be proposed that avoid a optimization on a min-max problem.
Acknowledgement: L. Jiang acknowledges the support of NSFC 12271408.